(-2^-2)^-1
1. Work inside parenthesis: (1/-2^2)^-1=(1/4)^-1
2. Clear exponent outside parenthesis: 4
This was marked as an incorrect answer. No correction given. Can you please help?

Answer by Theo(13342)

the answer should be -4
here's how it works.
(-2^-2)^-1 is equal to -2^(-2*-1) which is equal to -2^2) which is equal to -4
if you are working within the first parentheses first, the answer would be:
(-2^-2)^-1 = (-1/4)^-1 = -4
it's those negative signs that get you.
remember that:
-2^2 = (-1)*(2)^2 = -4
it is NOT equal to:
(-2)^2 = 4
i think that's where you went wrong.
handwritten solution is shown below:
$$$$$